9 Perkins Corporation is considering several investment proposals, as shown below: Investment Proposal A B C...

9

Perkins Corporation is considering several investment proposals, as shown below:

Investment Proposal
A B C D
Investment required $ 96,000 $ 120,000 $ 72,000 $ 90,000
Present value of future net cash flows $ 115,200 $ 180,000 $ 100,800 $ 168,000

If the project profitability index is used, the ranking of the projects from most to least profitable would be:

16

oetz Corporation has gathered the following data on a proposed investment project (Ignore income taxes.):

Investment required in equipment $ 40,000
Annual cash inflows $ 10,000
Salvage value of equipment $ 0
Life of the investment 15 years
Required rate of return 10 %

The company uses straight-line depreciation on all equipment. Assume cash flows occur uniformly throughout a year except for the initial investment.

Click here to view Exhibit 12B-1 and Exhibit 12B-2, to determine the appropriate discount factor(s) using the tables provided.

The internal rate of return of the investment is closest to:
Answer #1

9)

D,B,C,A

Profitability index = present value of future net cash flows / investment required

A = 115,200/96,000

= 1.2

B = 180,000/120,000

= 1.5

C = 100,800/72,000

= 1.4

D = 168,000/90000

= 1.87

16)

Annual cash inflows 10,000
Less: Depreciation 2,667 =40000/15
Annual net income 7,333
Annual net income 7,333
Divide by Investment 40,000
Simple rate of return 18.3%